Find the derivatives of this function


In(ye^x)=3


 

Grade:

2 Answers

arka mitra
12 years ago

ln(ye^x)=3

ye^x=e^3  (as per the property, logea=b means a=eb )

now, y=e^(3-x)

dy/dx=e^(3-x)d(3-x)/dx  (using chain rule)

:. dy/dx=-e^(3-x)

hope this helps....;)


Sanjeev Malik IIT BHU
12 years ago

ln(yex)=3

=>  lny + lnex =3 (using property)

=>  lny + x*lne =3

=> lny + x =3    (lne=1; using property-same baseand no.)

now diff;

dy/dx*1/y  +1 =0

so dy/dx= -y

 

 

please approve by clicking "yes" below